# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ IV.2 The conditions for enhanced-by-consistency SNR to occur
In general, we observe a non-monotonous behavior in both ${\tau_{\rm c}}$ and $\Delta{\mu_{\rm I}}$. The first one is the behavior that we described in Sec. IV.1. The non-monotonous behavior in $\Delta{\mu_{\rm I}}$ occurs since, for low input rate increment, the readout rate increment is negligible, while for too large increment ${\mu_{\rm I}}^{(+)}$ approaches $\hat{\mu_{\rm I}}$, and we loose the necessary condition for the SNR to increase.
@@
In Fig. 10 we see how the SNR contour plots change with the value of ${\mu_{\rm I}}^{(-)}$. For low enough ${\mu_{\rm I}}^{(-)}$, the signal increment is essentially given by ${r_{\rm R}}^{(+)}$, and the picture is as in Fig. 9: increasing ${\mu_{\rm I}}^{(-)}$ and $\Delta{\mu_{\rm I}}$ by the same quantity, leads to the same value of the SNR. The consequence is that increasing ${\mu_{\rm I}}^{(-)}$ has the same effect of a shift in the y-axis (Figs. 10-A,B). However, beyond a certain value of ${\mu_{\rm I}}^{(-)}$, ${r_{\rm R}}({\mu_{\rm I}}^{(-)})>{r_{\rm w}}({\mu_{\rm I}}^{(-)})$ is no longer negligible, and increasing ${\mu_{\rm I}}^{(-)}$ amounts to reduce $\Delta{r_{\rm R}}$, to increase ${v_{\rm R}}^{(-)}$ and, consequently, to a reduction of the SNR (Fig. 10-C, and Fig. 21 in Appendix L).
@@
Figure 10: Contour plot of the excess of SNR ($\delta_{\rm SNR}$) with respect to the absence of noise correlations, versus ${\tau_{\rm c}}$ and $\Delta{\mu_{\rm I}}$. Different panels correspond to different values of ${\mu_{\rm I}}^{(-)}=10,20,40$ (A, B, C, respectively). The theoretical SNR has been calculated according to the Poisson-Moreno (low $\alpha_{\rm M}$) and quenched-noise (large $\alpha_{\rm M}$) approximations. The thick purple line signals the threshold $\alpha_{\rm M}=0.75$ that we have used to select the approximation: above and at the left of the purple line, we use Poisson-Moreno, while below and at the right, we use the quenched-noise approximation. The dashed green line indicates the the parameter values for which ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$ (a necessary condition for the SNR to increase, assuming a negligible ${\mu_{\rm I}}^{(-)}$). The white dashed line indicates ${\tau_{\rm m}}$, while the dashed black line indicates the $\delta_{\rm SNR}=0$ contour. The rest of the model parameters are: ${\alpha_{\rm V}}=0.9$, ${\tau_{\rm m}}=0.05$, ${\sigma_{\rm V}}=1.104$, $w=0.45$.
@@
With similar arguments one can explain the qualitative dependence of the SNR on ${\tau_{\rm m}}$, $w$ and ${\sigma_{\rm c}}$ (see Appendix M).
@@ V Discussion
We formally analyzed a mechanism of enhanced-by-consistency readout proposed in [32], in which the readout is implemented as a single neuron with an output coincidence-detection nonlinearity, and studied it as a function of the interplay between the encoding and readout timescales. Previous work [32] analyzed this NERM by numerical integration, with ${N}=2$ encoding neurons. Here, we extend this work by providing an approximated analytic solution of a more general version of the model, which helps assessing the conditions under which the apparently paradoxical effect of lower encoding but higher transmission of neural information occurs. Previous works have addressed, with analytical approximations, the effect of temporal input correlations on the mean output rate of a LIF neuron [10, 16, 17]. Here, we provided alternative, approximated analytical expressions not only for the mean output firing rates (that actually generalize those in [17]), but also for the SNR of information transmission. This constitutes a key extension to evaluate the implications of correlated activity on neural information processing.
@@
We found that the non-linearity of the LIF activation function induces asymmetric fluctuations of the output firing rate, following symmetric input fluctuations. The quenched-noise approximation developed here, valid for large or moderate values of ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$, provides the conditions under which these asymmetric fluctuations allow information-limiting noise correlations to enhance the readout SNR. In particular, the following necessary conditions for the mean input current must be satisfied (see Sec. IV.2): the readout neuron must operate in a fluctuation-driven regime for low enough ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$, in agreement with the results of [32], with $\hat{\mu_{\rm I}}({v_{\rm c}},{\tau_{\rm m}})$ being a decreasing function of ${v_{\rm c}}$ and ${\tau_{\rm m}}$; the stimulus-specific difference in encoding firing rates ${\mu_{\rm I}}^{(+)}-{\mu_{\rm I}}^{(-)}$ is high enough (with ${\mu_{\rm I}}^{(-)}<\hat{\mu_{\rm I}}^{(-)}({v_{\rm c}},{\mu_{\rm I}})$, being $\hat{\mu_{\rm I}}^{(-)}({v_{\rm c}},{\tau_{\rm m}}$) a decreasing function of ${v_{\rm c}}$ and ${\tau_{\rm m}}$). These conditions on the input firing rates confirm previous conclusions that input correlations enhance readout information from a postsynaptic neuron when the readout integration time constant is short enough so that the average excitatory inputs received during an integration window are much smaller than the gap between the spiking threshold and resting potential of the readout neuron, so that output firing is driven by input fluctuations [32].
@@
Beyond the conditions on the encoding firing rates ${\mu_{\rm I}}^{(\pm)}$, our analytical solutions reveal necessary conditions on other variables, which were not considered in previous work [32] (see Appendix M). For instance, it reveals conditions on the amplitude of noise correlations, ${\sigma_{\rm c}}$: this quantity must be large enough for noise correlations to have an impact on the SNR numerator $\Delta{r_{\rm R}}$, but not too large for the SNR denominator ${v_{\rm R}}^{(s)}\right>_{s=\pm}$ to overcome the enhancement in the numerator (Fig. 9 and Sec. IV.2). The non-monotonic dependence of the readout SNR on ${\sigma_{\rm c}}^{2}=2{v_{\rm c}}{\tau_{\rm c}}$ and, hence, on ${\tau_{\rm c}}$ for fixed ${v_{\rm c}}$ (Fig. 8) can be clearly understood within the quenched-noise approximation. On the one hand, the larger ${\tau_{\rm c}}$, the larger the minimum value of the amplitude ${\sigma_{\rm c}}$ of noise correlations that can elicit a rate enhancement $\Delta{r_{\rm R}}$ (which depends on ${\sigma_{\rm c}}$ only through ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$). On the other hand, while $\Delta{r_{\rm R}}$ depends on the stationary variance ${v_{\rm c}}$ only, the readout variance ${v_{\rm R}}$ increases with both ${v_{\rm c}}$ and ${\tau_{\rm c}}$ (and faster with ${v_{\rm c}}$, cf. Eq. (13a)). The dependence of the mean readout rate ${r_{\rm R}}$ on the stationary variance ${v_{\rm c}}$ follows from the quenched-noise approximation ansatz itself (Eq. 9). In contrast, the readout rate variance ${v_{\rm R}}$ increases with ${v_{\rm c}}$ and ${\tau_{\rm c}}$, because it is the integration, Eq. (11), of the temporal correlation function (which, in the quenched-noise approximation, decays as an exponential with timescale ${\tau_{\rm c}}$ and amplitude $Q({v_{\rm c}})$, see Eqs. (12) and (72) in Appendix H). As a result, the enhancement of the readout SNR occurs over a wide range of timescales ${\tau_{\rm c}}$ (or ${\sigma_{\rm c}}$ for fixed ${v_{\rm c}}$), while the SNR decreases at sufficiently large ${\tau_{\rm c}}$.
@@
The additional necessary conditions derived here have potential implications for the interpretation of biological findings and functions. We have found that information-limiting correlations are useful to enhance information transmission only for intermediate values of the correlation strength ${\sigma_{\rm c}}$. Across experiments and brain areas, correlation between neurons are reported consistently to be non-zero, but always of small or moderate value [23, 8]. Our result suggests that non-null but moderate values may be useful to design neural codes that trade off effectively the information-limiting and information-enhancing effect of correlations and achieve good levels of information transmission. We have also found that the SNR enhancement may occur for a wide range of time constants of input correlations ${\tau_{\rm c}}$, even at values of ${\tau_{\rm c}}$ much larger than the readout membrane time constant ${\tau_{\rm m}}$. This implies that neural circuits may be able to enjoy the benefits of information-encoding-limiting yet information-transmission-enriching correlations even when encoding information with activity correlated over long times. This is important because long encoding timescales are useful to combine temporally separate inputs and accumulate consistently information over time, or to implement or compute a goal-oriented navigation plan [34].
@@
Acknowledgements
We acknowledge Sandro Azaele for pointing out references [56, 57, 58, 59]. M. I. is supported by: the European Union -Horizon 2020 Program under the scheme INFRAIA-01-2018-2019 –Integrating Activities for Advanced Communities, Grant Agreement n.871042, “SoBigData++: European Integrated Infrastructure for Social Mining and Big Data Analytics” (http://www.sobigdata.eu); by the European Union - NextGenerationEU - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR), project ‘SoBigData.it - Strengthening the Italian RI for Social Mining and Big Data Analytics’ - Grant IR0000013 (n. 3264, 28/12/2021) (https://pnrr.sobigdata.it/); by the Spanish Ministry and Agencia Estatal de Investigación (AEI) through Project I+D+i Ref. No. PID2020-113681GB-I00, financed by MICIN/AEI/10.13039/501100011033 and FEDER, “A way to make Europe”.
@@ A.2 Table of symbols and acronyms
$\rho_{i}(t)$ $=\sum_{k>0}\delta(t-t_{k}^{(i)})$ density of spikes of encoding neuron $i$ (single realisation)
$\rho(t)$ $=\sum_{j}\rho_{j}(t)$ density of encoding spikes in a single realisation
$n({T_{\rm c}})$ $\int_{0}^{T_{\rm c}}{\rm d}t\rho(t)$ number of encoding spikes in a single realisation
@@
${n_{\rm R}}({T_{\rm c}})$ $\int_{0}^{T_{\rm c}}{\rm d}t{\varrho}(t)$ number of readout spikes in a single realisation
@@
$\hat{\mu_{\rm I}}$ upper bound of ${\mu_{\rm I}}^{(+)}$ for the SNR to increase (see IV.2)
$\hat{\mu_{\rm I}}^{(-)}$ lower bound of ${\mu_{\rm I}}^{(-)}$ for the SNR to increase (see IV.2)

### v2

@@ IV.3 The conditions for enhanced-by-consistency SNR to occur
Dependence on other variables. These considerations are enough to understand the qualitative dependence of the SNR with respect to the rest of the LIF parameters. We exemplify such dependencies in Fig. 11 and in Figs. 32,33,34 in Appendix O. In general, we observe a non-monotonous behavior in both ${\tau_{\rm c}}$ and $\Delta{\mu_{\rm I}}$. The first one is the behavior that we described in Sec. IV.2. The non-monotonous behavior in $\Delta{\mu_{\rm I}}$ occurs since, for low input rate increment, the readout rate increment is negligible, while for too large increment ${\mu_{\rm I}}^{(+)}$ approaches $\hat{\mu_{\rm I}}$, and we loose the necessary condition for the SNR to increase.
@@
In Fig. 11 we see how the SNR contour plots change with the value of ${\mu_{\rm I}}^{(-)}$. For low enough ${\mu_{\rm I}}^{(-)}$, the signal increment is essentially given by ${r_{\rm R}}^{(+)}$, and the picture is as in Fig. 10: increasing ${\mu_{\rm I}}^{(-)}$ and $\Delta{\mu_{\rm I}}$ by the same quantity, leads to the same value of the SNR. The consequence is that increasing ${\mu_{\rm I}}^{(-)}$ has the same effect of a shift in the y-axis (Figs. 11-A,B). However, beyond a certain value of ${\mu_{\rm I}}^{(-)}$, ${r_{\rm R}}({\mu_{\rm I}}^{(-)})>{r_{\rm w}}({\mu_{\rm I}}^{(-)})$ is no longer negligible, and increasing ${\mu_{\rm I}}^{(-)}$ amounts to reduce $\Delta{r_{\rm R}}$, to increase ${v_{\rm R}}^{(-)}$ and, consequently, to a reduction of the SNR (Fig. 11-C, and Fig. 30 in Appendix N).
@@
Figure 11: Contour plot of the excess of SNR ($\delta_{\rm SNR}$) with respect to the absence of noise correlations, versus ${\tau_{\rm c}}$ and $\Delta{\mu_{\rm I}}$. Different panels correspond to different values of ${\mu_{\rm I}}^{(-)}=10,20,40$ (A, B, C, respectively). The theoretical SNR has been calculated according to the Poisson-Moreno (low $\alpha_{\rm M}$) and quenched-noise (large $\alpha_{\rm M}$) approximations. The thick purple line signals the threshold $\alpha_{\rm M}=0.75$ that we have used to select the approximation: above and at the left of the purple line, we use Poisson-Moreno, while below and at the right, we use the quenched-noise approximation. The dashed green line indicates the the parameter values for which ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$ (a necessary condition for the SNR to increase, assuming a negligible ${\mu_{\rm I}}^{(-)}$). The white dashed line indicates ${\tau_{\rm m}}$, while the dashed black line indicates the $\delta_{\rm SNR}=0$ contour. We remind that ${\sigma_{\rm c}}$ varies with ${\tau_{\rm c}}$ according to Eq. (6c). The rest of the model parameters are: ${\tau_{\rm m}}=0.05$, ${\alpha_{\rm V}}=0.9$, ${N}=20$, ${\sigma_{\rm V}}=1.104$, $w=0.45$.
@@
With similar arguments one can explain the qualitative dependence of the SNR on ${\tau_{\rm m}}$, $w$ and ${\sigma_{\rm c}}$ (see Appendix O).
IV.4 Further analyses of the SNR in the NERM
Dependence on ${N}$. For completeness, we have carried out two further analyses regarding the behavior of the SNR in the NERM in the diffusion approximation. The first one, in Appendix D.3 addresses the behavior for large ${N}$, whenever the single encoding neuron parameters $\mu_{i},{\sigma_{\rm V}}_{i}$ are naturally rescaled with ${N}$ as $\mu_{i}=\mu_{i}^{({N}_{0})}{N}_{0}/{N}$, ${\sigma_{\rm V}}={\sigma_{\rm V}}^{({N}_{0})}{N}_{0}/{N}$. In this scenario, the NERM in Eq. (5) is expected to become independent of $N$, as the only relevant parameters ${\mu_{\rm I}},{\sigma_{\rm I}},{\sigma_{\rm c}}$ governing the encoding signal in the NERM take the form (large $N$, see Eq. 6): ${\mu_{\rm I}}=w\mu_{i}^{({N}_{0})}$, ${\sigma_{\rm I}}^{2}=w^{2}\mu_{i}^{({N}_{0})}$, ${\sigma_{\rm c}}\simeq w{\sigma_{\rm V}}^{({N}_{0})}{\alpha_{\rm V}}$. The numerical simulations in Appendix D.3 are consistent with this picture.
Varying ${\sigma_{\rm c}}$ at constant ${\tau_{\rm c}}$. The second complementary analysis (see Appendix F) regards the behavior of the SNR in the NERM for fixed ${\tau_{\rm c}}$ and varying ${\sigma_{\rm V}}$ (varying, equivalently, ${\sigma_{\rm c}}=w{\sigma_{\rm V}}N(1+(N-1){\alpha_{\rm V}}^{2})^{1/2}$). Besides the numerical underestimation of the rate variance induced by the finite-NIS effect, the theory provides us with a correct picture of the non monotonous behavior of the SNR with ${\sigma_{\rm c}}$, as it did in the above analysis for constant ${v_{\rm c}}$ and varying ${\tau_{\rm c}}$.
@@ V Discussion
We formally analyzed a mechanism of enhanced-by-consistency readout[32], in which the readout is implemented as a single neuron with an output coincidence-detection nonlinearity, and studied it as a function of the interplay between the encoding and readout timescales. Previous work [32] analyzed this NERM by numerical integration, with ${N}=2$ encoding neurons. Here, we extend this work by providing an approximated analytic solution of a more general version of the model, which helps assess the conditions under which the apparently paradoxical effect of lower encoding but higher transmission of neural information occurs. Previous works have addressed, with analytical approximations, the effect of temporal input correlations on the mean output rate of a LIF neuron [10, 16, 17]. Here, we provided alternative, approximated analytical expressions not only for the mean output firing rates (that actually generalize those in [17]), but also for the SNR of information transmission. This constitutes a key extension to evaluate the implications of correlated activity on neural information processing.
@@
We found that the non-linearity of the LIF activation function induces asymmetric fluctuations of the output firing rate, following symmetric input fluctuations. The quenched-noise approximation developed here, valid for large or moderate values of ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$ and for low stationary noise variance ${v_{\rm c}}$, provides us with the conditions under which these asymmetric fluctuations allow information-limiting noise correlations to enhance the readout SNR. In particular, the following necessary conditions for the mean input current must be satisfied (see Sec. IV.3): the readout neuron must operate in a fluctuation-driven regime for low enough ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$, in agreement with the results of [32], with $\hat{\mu_{\rm I}}({v_{\rm c}},{\tau_{\rm m}})$ being a decreasing function of ${v_{\rm c}}$ and ${\tau_{\rm m}}$; the stimulus-specific difference in encoding firing rates ${\mu_{\rm I}}^{(+)}-{\mu_{\rm I}}^{(-)}$ is high enough (with ${\mu_{\rm I}}^{(-)}<\hat{\mu_{\rm I}}^{(-)}({v_{\rm c}},{\mu_{\rm I}})$, being $\hat{\mu_{\rm I}}^{(-)}({v_{\rm c}},{\tau_{\rm m}}$) a decreasing function of ${v_{\rm c}}$ and ${\tau_{\rm m}}$). These conditions on the input firing rates confirm previous conclusions that input correlations enhance readout information from a postsynaptic neuron when the readout integration time constant is short enough so that the average excitatory inputs received during an integration window are much smaller than the gap between the spiking threshold and resting potential of the readout neuron, so that output firing is driven by input fluctuations [32].
@@
Beyond the conditions on the encoding firing rates ${\mu_{\rm I}}^{(\pm)}$, our analytical solutions reveal necessary conditions on other variables, which were not considered in previous work [32]. For instance, it reveals conditions on the amplitude of noise correlations, ${\sigma_{\rm c}}$: this quantity must be large enough for noise correlations to have an impact on the SNR numerator $\Delta{r_{\rm R}}$, but not too large for the SNR denominator ${v_{\rm R}}^{(s)}\right>_{s=\pm}$ to overcome the enhancement in the numerator (Fig. 10 and Sec. IV.3). The non-monotonic dependence of the readout SNR on ${\sigma_{\rm c}}^{2}=2{v_{\rm c}}{\tau_{\rm c}}$ and, hence, on ${\tau_{\rm c}}$ for fixed ${v_{\rm c}}$ (Fig. 9) can be clearly understood within the quenched-noise approximation. On the one hand, the larger ${\tau_{\rm c}}$, the larger the minimum value of the amplitude ${\sigma_{\rm c}}$ of noise correlations that can elicit a rate enhancement $\Delta{r_{\rm R}}$ (which depends on ${\sigma_{\rm c}}$ only through ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$). On the other hand, while $\Delta{r_{\rm R}}$ depends on the stationary variance ${v_{\rm c}}$ only, the readout variance ${v_{\rm R}}$ increases with both ${v_{\rm c}}$ and ${\tau_{\rm c}}$ (and faster with ${v_{\rm c}}$, cf. Eq. (13a)). The dependence of the mean readout rate ${r_{\rm R}}$ on the stationary variance ${v_{\rm c}}$ follows from the quenched-noise approximation ansatz itself (Eq. 9). In contrast, the readout rate variance ${v_{\rm R}}$ increases with ${v_{\rm c}}$ and ${\tau_{\rm c}}$, because it is the integration, Eq. (11), of the temporal correlation function (which, in the quenched-noise approximation, decays as an exponential with timescale ${\tau_{\rm c}}$ and amplitude $Q({v_{\rm c}})$, see Eqs. (12) and (110) in Appendix I). As a result, the enhancement of the readout SNR can occur over a wide range of timescales ${\tau_{\rm c}}$ (or ${\sigma_{\rm c}}$ for fixed ${v_{\rm c}}$), while it decreases at sufficiently large noise amplitude ${\sigma_{\rm c}}$.
@@
It is of interest to consider the neuroscientific implications of the conditions and parameter regions we found to be able to support an enhanced-by-consistency information transmission regime. The first consideration regards how the numerical range of parameters in which this is possible compares with real biological parameters. Figure 11 explores this regime as a function of two parameters that have been extensively measured in the brain: the time constant of correlations $\tau_{c}$ and the firing rate differences between the different stimuli to be detected. The correlation time constant increases systematically when moving from sensory to higher-order areas [53], and it ranges from one or few milliseconds for the retina or sub-cortical structures [54, 55, 56], to few tens of milliseconds for cortical sensory areas, and hundreds of milliseconds for association or decision-making areas [34, 57, 58]. The firing rate differences between stimuli to be discriminated can vary from fractions of one Hz (spikes/s) to few tens of Hz [59, 60, 61, 62, 63, 64]. Our results (Fig. 11) show an enhanced-by-consistency regime for rate differences spanning few tens of Hz and correlation time constants r $\tau_{c}$ larger than the membrane time constant $\tau_{m}$ (which is typically in the range of 5–20 ms) and reaching several tens to few hundreds of millisecond, well within the range of sensory and association areas, suggesting that the enhanced-by-consistency regime can be sustained in these brain areas. The second consideration is that the additional necessary conditions derived here have potential implications for the neural biological functions. We have found that information-limiting correlations are useful to enhance information transmission only for intermediate values of the correlation strength ${\sigma_{\rm c}}$. Across experiments and brain areas, correlation between neurons are reported consistently to be non-zero, but always of small or moderate value [23, 8]. Our result suggests that non-null but moderate values may be useful to design neural codes that trade off effectively the information-limiting and information-enhancing effect of correlations and achieve good levels of information transmission. We have also found that the SNR enhancement may occur for a wide range of time constants of input correlations ${\tau_{\rm c}}$, even at values of ${\tau_{\rm c}}$ much larger than the readout membrane time constant ${\tau_{\rm m}}$. This implies that neural circuits may be able to enjoy the benefits of information-encoding-limiting yet information-transmission-enriching correlations even when encoding information with activity correlated over long times. This is important because long encoding timescales are useful to combine temporally separate inputs and accumulate consistently information over time, or to implement or compute a goal-oriented navigation plan [34].
@@
The function of populations of sensory neurons includes encoding the time course of time-varying stimuli, beyond the encoding of the identity of static stimuli as considered in our study. It would be important in future studies to extend our work to consider how noise correlations and readout nonlinearities impact the encoding and readout of information about the time course of time-varying stimuli. Indeed, an empirical study [66] suggested that noise correlations limit the temporal precision of the information that the time course of the average firing in a neural population can convey about the time course of a stimulus. Other studies have proposed that readout nonlinearities can enhance the detection of very rapid stimulus variations in a way that is robust to the detrimental encoding effects of noise correlations [67]. While studying the time-varying encoding and readout is beyond the scope of the present article, we speculate that extending our formalism to this problem will give insights about the precise conditions in which the temporal parameters of stimulus dynamics, noise correlations and readout nonlinearities allow for a beneficial effect of time-dependent noise correlations on the transmission of information about dynamic stimuli.
@@
Acknowledgments
We acknowledge Sandro Azaele for pointing out references [70, 71, 72, 73]. M. I. is supported by: the European Union under the scheme HORIZON-INFRA-2021-DEV-02-01 – Preparatory phase of new ESFRI research infrastructure projects, Grant Agreement n.101079043, “SoBigData RI PPP: SoBigData RI Preparatory Phase Project”; by the project “Reconstruction, Resilience and Recovery of Socio-Economic Networks” RECON-NET EP_FAIR_005 - PE0000013 “FAIR” - PNRR M4C2 Investment 1.3, financed by the European Union – NextGenerationEU; by the European Union - Horizon 2020 Program under the scheme ‘INFRAIA-01-2018-2019 - Integrating Activities for Advanced Communities’, Grant Agreement n.871042, ‘SoBigData++: European Integrated Infrastructure for Social Mining and Big Data Analytics’ (http://www.sobigdata.eu); Grant No. PID2023-149174NB-I00 financed by MICIU/AEI/10.13039/501100011033 and EDRF/EU funds. M. I. also benefits from the IMT computing resources funded by the DM, Italy n. 289, 25-03-2021, CUP D67G22000130001, through the project “Scientific computing for natural sciences, social sciences, and applications: methodological and technological development”. S. P. is supported by the the German Federal Ministry of Education and Research, grant BMBF 01GQ2404.
@@ A.2 Table of symbols and acronyms
$\rho_{i}(t)$ $=\sum_{k>0}\delta(t-t_{k}^{(i)})$ density of spikes of encoding neuron $i$ (single realization)
$\rho(t)$ $=\sum_{j}\rho_{j}(t)$ density of encoding spikes in a single realization
$n({T_{\rm c}})$ $\int_{0}^{T_{\rm c}}{\rm d}t\rho(t)$ number of encoding spikes in a single realization
@@
${n_{\rm R}}({T_{\rm c}})$ $\int_{0}^{T_{\rm c}}{\rm d}t{\varrho}(t)$ number of readout spikes in a single realization
@@
$\hat{\mu_{\rm I}}$ upper bound of ${\mu_{\rm I}}^{(+)}$ for the SNR to increase (see IV.3)
$\hat{\mu_{\rm I}}^{(-)}$ lower bound of ${\mu_{\rm I}}^{(-)}$ for the SNR to increase (see IV.3)